# From Challenges and Pitfalls to Recommendations and Opportunities: Implementing Federated Learning in Healthcare

### v1

@@ 5.4.2 Open Domain Problem
Studies have shown that even slight differences in devices or acquisition protocols can result in a significant distribution shift, thereby reducing the model’s effectiveness when applied to new, unseen datasets. This issue is particularly acute in healthcare applications like diabetic retinopathy screening in fundus images, where the diversity of cameras and settings across different institutions can lead to poor model performance on external data. 31 addressed this challenge by introducing a frequency-based domain generalization approach in FL. They enabled privacy preserving exchange of distribution information across clients through continuous frequency space interpolation and designed a boundary-oriented episodic learning scheme to expose local training to domain shifts and enhance model generalizability in ambiguous boundary regions. However, the proposed method can be impractical for real-world applications due to its reliance on extensive network bandwidth and computational resources required for Fourier transform computation in frequency space interpolation.
@@
  • Domain Generalization Techniques: These techniques aim to create models that are robust to distribution shifts by learning domain-invariant features. Methods such as data augmentation, adversarial training, and meta-learning have been explored to improve the generalization capabilities of FL models to unseen domains.
  • Transfer Learning and Fine-Tuning: Apply transfer learning and fine-tuning on small amounts of unlabeled data from the open domain can help adapt the federated model to new environments 77. Self-supervised learning techniques like contrastive learning can further facilitate this process 37.
  • Ensemble Learning: Combine multiple models trained on different clients via sophisticated ensemble methods can enhance the robustness of the final prediction, making it more generalizable to open domains 151.
  • Data Synthesis and Simulation: Generate synthetic data that mimics the characteristics of potential open domains can be used to pretrain or fine-tune the federated model, improving its generalization to unseen environments 31.
@@ 5.4.3 Burdens in Communication
Communication is a significant bottleneck in the implementation of FL in healthcare. In FL, each client needs to frequently communicate with the central server. This communication can be orders of magnitude slower than local computation due to constraints such as bandwidth, latency, and power 95.
The communication bottleneck in FL arises from several factors. First, the number of clients involved in an FL system can be very large (e.g., wearables and IoMT), leading to significant communication overhead. Each communication round requires sending model updates between the clients and the central server, which can be expensive in terms of time and resources 95, 4. Second, ensuring data privacy and security in FL is crucial, especially in sensitive domains like healthcare. The need for encryption and secure communication protocols adds to the computational and communication overhead. Encrypting model updates can significantly increase the size of the data being transmitted, further straining the communication channels and requiring more sophisticated algorithms to balance privacy and efficiency 4. Third, as FL tasks become more complex, the size of the models involved increases. Modern large-scale models, such as large language models (LLMs) and foundation models (FMs), can have billions of parameters, resulting in model sizes that require significant bandwidth to transmit. This issue is exacerbated when using standard communication protocols like gRPC, which have size limits on single messages (e.g., 2 GB). Typical LLMs and FMs can exceed these limits, necessitating the model to be split into smaller chunks for transmission, adding additional overhead and complexity to the system 158.
Included studies primarily concentrated on enhancing communication encryption techniques, with the aim of either reducing the volume of data exchanged 63, 104 or minimizing the number of communication rounds 54. Additionally, several studies explored methods for achieving fully decentralized communication without central server 55, 159. Improvements in encryption often involved methods for securely sharing secret keys among clients 83, 81, encryption mechanisms for safeguarding exchanged data 82, 64, and techniques for perturbing model outputs at each client using a secret key 160. To decrease the amount of data transmitted, some studies proposed transferring only a subset of model parameters 141, 161, 95 or employing strategies like compressing 97, masking 61, and quantizing gradients 95 or model outputs before exchange 27. Reducing the number of communication rounds was addressed through model design 91, 162, 98, aggregating updates based on elapsed time instead of epochs 97, 107, and evaluating the potential benefit of an update before communication 81. Other studies focused on detecting attacks during communication 163, developing authentication systems for clients 46, and improving client management systems 45, 64.
@@
  • Model Optimization: Utilize parameter-efficient, lightweight models such as MobileNet and SqueezeNet to reduce computation and memory overhead. Explore model pruning 164 and low-rank adaptation (LoRA) 165 techniques to decrease model size or regularize model weights without significantly impacting accuracy. Consider split learning 101, 25 to share intermediate activations or a subset of the model instead of the full model, thus reducing communication costs.
@@
  • Reduce Communication Rounds: Minimize the number of communication rounds by adopting methods such as One-Shot FL 166, which requires only a single round of communication to achieve effective model training.
@@
  • Knowledge Distillation: Use knowledge distillation techniques to transmit only essential distilled knowledge, such as logits from the final layer, rather than the entire model 152, 166, 50. This approach can reduce communication overhead while maintaining model performance.
  • Data Distillation: Apply data distillation to generate synthetic data summaries or distilled representations of the original datasets or updates. Clients can then share smaller, more concise updates with the central server, enhancing communication efficiency 167, 168.
@@ 5.4.4 Plain Convergence Analysis
Federated optimization in healthcare aims to adapt models to local data distributions while integrating global information. The inherent heterogeneity across hospitals and institutions often leads to instability and slow convergence in federated training 154. However, comprehensive convergence analysis is frequently lacking in current studies.
@@
Only a handful of studies 159, 61, 95, 104 have focused on the convergence of FL in healthcare settings. These studies typically relied on Stochastic Gradient Descent (SGD) as the foundational optimization method due to its effectiveness in smooth optimization problems, under assumptions such as the existence of lower bounds, Lipschitz smoothness, and bounded variance 159, 61. However, SGD-based FL algorithms often struggle with nonsmooth optimization problems, which are common in healthcare data due to irregularities in data distributions and the presence of outliers.
@@
  • Theoretical Convergence Guarantees: Establish rigorous theoretical methods that provide convergence guarantees for FL in healthcare. This includes deriving bounds on convergence rates and understanding the conditions under which proposed algorithms perform optimally, particularly in non-IID data settings.
@@
  • Advanced Optimization: Develop optimization algorithms tailored to the specific challenges of healthcare data, such as non-convexity, non-smoothness, and heterogeneity. Techniques like adaptive learning rates, momentum methods, and second-order methods could be explored to enhance convergence rates and stability in FL scenarios.
@@ 5.4.5 Temporal Dynamics and Revoke Issues
Healthcare data’s inherent time dependence is critical, especially for diseases with distinct progression or treatment timelines, such as cancer and chronic conditions like diabetes. However, included studies often overlooked these temporal dynamics when partitioning data, potentially leading to models that inaccurately reflect disease progression. For example, COVID-19 characteristics, such as ground-glass opacities in lung CT scans, evolve with the disease’s progression 169. Ignoring such temporal dynamics can result in models that are overfitted to specific stages of a disease and unable to generalize across different phases 170, 171, 172. Furthermore, the dynamic nature of data at each participating hospital or institution complicates the situation. Hospitals continuously acquire new data and may also remove or modify existing data due to errors or other factors. This dynamic data environment requires FL models to adapt without frequent retraining, as new data might cause model drift, while data removal can leave critical gaps in the model’s understanding, particularly if the removed data represents rare or critical cases.
Another critical but overlooked issue in FL is data revocation, which becomes necessary when specific data must be withdrawn due to privacy concerns, regulatory requirements, patient requests, or participant requests. Current FL setups, designed for iterative data aggregation, struggle with “unlearning” specific contributions without requiring complete model retraining. Emerging research highlights the need for mechanisms that allow for efficient data revocation without compromising the integrity of the model. For instance, methods have been proposed to facilitate client “unlearning” in FL, enabling the removal of data contributions from specific clients without significant degradation in model performance 173. This is especially important in healthcare, where patient consent might be withdrawn, or new privacy laws might mandate the deletion of certain data.
@@
  • Data Revocation: Develop and integrate efficient data revocation techniques in FL frameworks. Approaches such as machine unlearning 173 can be refined to allow the removal of specific data contributions while minimizing the impact on overall model performance. This will be critical for maintaining compliance with privacy laws and upholding patient rights in healthcare applications.
@@ 5.4.6 Synchronization Issues
In Synchronous FL, all clients must complete their training and send their updates before the global model aggregation, which is straightforward but can be inefficient. This approach works well in environments where data is immediately available, such as in a centralized hospital picture archiving and communication system (PACS). However, in real-world scenarios, where data acquisition might be delayed due to network issues or the unavailability of input/output devices, synchronous updates can result in significant idle times and resource underutilization. Additionally, clients in an FL network, particularly smaller healthcare entities (e.g., wearables), may not be active during every communication round, further delaying the global model update and potentially degrading overall system performance.
Asynchronous FL, on the other hand, allows clients to send updates independently, without waiting for other clients to finish their training. This approach is more flexible and can accommodate variations in client availability and computational power, thereby improving the overall efficiency of the FL process. However, asynchronous updates can introduce challenges related to the consistency of the global model, as updates from slower or less reliable clients may arrive out of sync with the rest of the system.
@@
Model Inversion 174 Reconstruct the actual samples (e.g., patient medical images or genetic data) from the model or updates Potential leakage of original patient data High Data privacy, patient confidentiality Anomaly detection in model outputs & updates DP
Membership Inference 175 Determining if a specific patient’s record was part of the model’s training set by analyzing its outputs Compromises patient confidentiality, leading to potential unauthorized access to sensitive health information Medium Data privacy, patient confidentiality Monitoring for unusual model behaviour, particularly overconfident predictions Regularization techniques, DP, and robust model validation
Data Attribute Inference 176 Infer individual sample’s attributes (like race, gender, age) or gain aggregate statistical insights about the entire training set from model parameters and updates Leakage of patient privacy Medium Data privacy, patient confidentiality, communication security Gradient analysis, privacy audits Secure aggregation, DP
Data Poisoning 177 Malicious participants alter local data or labels to degrade the global model’s performance Significant drop in model performance, potentially misleading patient diagnoses Medium Model accuracy, patient safety Monitoring for abnormal model behaviour, statistical checks Robust aggregation methods, anomaly detection in model updates
Model Poisoning 72 Malicious participants upload tampered model parameters to manipulate the global model Global model may be intentionally corrupted, affecting its performance on real data High Model integrity, patient safety Anomaly detection in model updates Robust aggregation, DP
Denial of Service 73 Disrupt the FL system by overwhelming it with requests or blocking normal data flow Training process delays or interruptions, affecting time-sensitive medical applications Medium System availability and security, training efficiency Monitoring network traffic for anomalies Rate limiting, robust network design, redundancy mechanisms
HE Attack 64 Decrypt encrypted model updates to access sensitive information Encryption key leakage may lead to a breakdown in data protection, compromising patient privacy High Data privacy, model integrity Encryption integrity checks, key management audits Secure key management, SMPC
User Withdrawal Attack 44 Participants withdraw from training, but their prior updates still affect the global model Updates from withdrawn users may contain errors or malicious data, degrading the global model Low Model integrity Tracking user participation and contribution consistency Periodic re-evaluation of model contributions, machine unlearning
@@ 5.5 Privacy and Security
Two critical privacy and security issues exist in current studies. First, it is concerning that 62% of the included studies did not encrypt model updates during communication. This lack of encryption leaves FL system vulnerable to interception, posing significant security risks. Second, statistical information such as sample sizes and distributions were often shared alongside model updates, particularly in FedAvg-based methods 106, 108. This practice can expose participants with large datasets to targeted attacks if an adversary intercepts the communication or compromises the aggregator 158.
Despite the advantage of FL in healthcare without directly exchanging or sharing local data, it is not immune to privacy and security risks. Adversaries can analyze changes in model updates over time to infer sensitive information or exploit system vulnerabilities to conduct targeted attacks using techniques like model inversion 178, membership inference 179, and poisoning 177. Additionally, clients may unintentionally reveal private data during the FL process. This can happen when the client memorizes previous model and gradient updates, leading to the leakage of sensitive information 65, 16. Furthermore, methods involving the sharing of a few data samples for augmentation or disclosing local data distributions during knowledge transfer can also result in privacy breaches 99, 50. These adversarial and unintentional exposures undermine the privacy and security of the FL process, necessitating robust countermeasures and continuous vigilance to safeguard the integrity and confidentiality of the FL process.
@@
Cryptography encompass a variety of methods, with homomorphic encryption (HE) 180 and secure multi-party computation (SMPC) 181 being among the most popular. HE enables computations directly on encrypted data without the need for decryption. This ensures that data remains encrypted throughout processing, storage, and transmission, thus providing robust data security and privacy. HE allows operations on ciphertexts, with the results, once decrypted, accurately reflecting the outcomes of operations performed on the original plaintext data. HE is particularly valuable in FL due to its ability to safeguard data privacy during computation. Recent research has highlighted HE’s effectiveness in various healthcare applications, including oncology and medical genetics. For instance, 69 demonstrated HE’s potential for truly private federated evaluations, and 77 successfully utilized HE for model aggregation in FL. While SMPC enables multiple parties to collaboratively compute a function over their combined data while keeping each party’s data private. Each participant holds a piece of encrypted or encoded data, and the computation is designed so that no party can access the others’ data or infer anything beyond the final result. SMPC ensures the confidentiality of both input data and computation results, making it robust against adversarial attacks and suitable for scenarios with multiple untrusted parties. It is increasingly used in healthcare and other sensitive applications to enhance privacy. For instance, research shows that SMPC can improve privacy in FL with medical datasets by addressing risks related to malicious models and enhancing the confidentiality of model aggregation 47, 72.
@@
Nevertheless, HE typically involves high storage and computational overheads. Encrypted data requires significant processing power, and the complexity of HE schemes can introduce a single point of failure, where a single server manages all encrypted data 69. Additionally, managing encryption keys securely is crucial, as any compromise in key management can jeopardize the entire system’s security 182. Also, SMPC often incurs high computational and communication overhead. The process of encrypting, encoding, and splitting data can be computationally intensive. Moreover, coordinating the computations across multiple parties requires substantial communication resources, which can become a bottleneck as the number of participants grows 183.
@@
Perturbation techniques, with differential privacy (DP) 184 being the most prominent, are crucial for protecting sensitive healthcare data. DP quantifies the risk of exposing individual data points and ensures that the inclusion or exclusion of any single data point has minimal impact on the model’s output by introducing randomness into the model’s results. In FL, DP is typically implemented by adding noise to the local updates or gradients before they are aggregated. This noise is designed to mask the contributions of individual data points, thus making it difficult to infer any single data point from the model’s outputs. DP provides a quantifiable measure of privacy through parameters such as $\epsilon$ and $\delta$, allowing for a clear understanding of the trade-off between privacy and model utility. Recent studies in healthcare have shown that FL with DP can achieve comparable accuracy to non-DP models in specific tasks, with a performance gap of less than 5%, proving DP’s efficacy with minimal accuracy compromise 106, 108, 158.
@@
However, the main challenge with DP is balancing the privacy budget with model performance. Adding noise typically reduces model accuracy by obscuring data patterns, which is a critical consideration in DP implementation 65. Managing the privacy budget $\epsilon$ involves a trade-off: a smaller $\epsilon$ enhances privacy but may reduce performance, and a larger $\epsilon$ offers less privacy protection. Effective management of this trade-off is essential, requiring careful attention to privacy needs and performance goals 128. Implementing DP also introduces additional computational overhead due to the noise addition and gradient clipping processes. This overhead can affect the scalability and efficiency of FL systems, particularly in scenarios with large numbers of clients or data 185.
@@
Beyond cryptography and perturbation methods, blockchain 23 and swarm learning (SL) 68, 24 have gained lots of attention for enhancing privacy and security in healthcare. Blockchain is increasingly being integrated with FL to address data security and trust issues by replacing the central server with a decentralized privacy protocol. The key advantage of blockchain lies in its distributed ledger system, which ensures that data is securely shared and maintained across all clients without relying on a central authority. This decentralization reduces the risk of single points of failure and enhances data integrity. Furthermore, blockchain’s verification schemes play a crucial role in the FL process, helping to detect and mitigate threats such as poisoning attacks. For instance, a blockchain-based FL framework developed by 75 combined DP and gradient-verification protocols to enhance security in IoMT devices, significantly reducing the success rate of poisoning attacks in tasks like diabetes prediction. Another approach by 74 utilized blockchain alongside an intrusion detection system to monitor and prevent malicious activities during model training, further securing federated healthcare networks. However, while blockchain offers robust security features, its integration with FL is often challenged by high communication and computational costs, which can limit its scalability and efficiency. Swarm Learning takes a different approach by decentralizing not just the privacy protocol but the entire model training process. In SL, there is no central aggregator, instead, decentralized hardware infrastructures work together to securely onboard clients and collaboratively generate a global model. This decentralized approach enhances the network’s resilience against attacks and is particularly effective in scenarios where data is non-IID, such as in the prediction of conditions like COVID-19 and leukemia 68, 24. However, SL’s reliance on peer-to-peer communication can introduce latency issues, which may slow down the training process, particularly in environments with varying network conditions.

### v2

@@ 5.4.2 Open Domain Problem
Studies have shown that even slight differences in devices or acquisition protocols can result in a significant distribution shift, thereby reducing the model’s effectiveness when applied to new, unseen datasets. This issue is particularly acute in healthcare applications like diabetic retinopathy screening in fundus images, where the diversity of cameras and settings across different institutions can lead to poor model performance on external data. 33 addressed this challenge by introducing a frequency-based domain generalization approach in FL. They enabled privacy preserving exchange of distribution information across clients through continuous frequency space interpolation and designed a boundary-oriented episodic learning scheme to expose local training to domain shifts and enhance model generalizability in ambiguous boundary regions. However, the proposed method can be impractical for real-world applications due to its reliance on extensive network bandwidth and computational resources required for Fourier transform computation in frequency space interpolation.
@@
  • Domain Generalization Techniques: These techniques aim to create models that are robust to distribution shifts by learning domain-invariant features. Methods such as data augmentation 163, which generates diverse augmented features of client data to improve model robustness, adversarial training 164, where domain-invariant representations are learned by minimizing domain discrimination, and meta-learning 165, which optimizes for rapid adaptation to new domains, have been explored to improve the generalization capabilities of FL models to unseen domains.
  • Transfer Learning and Fine-Tuning: Apply transfer learning and fine-tuning on small amounts of unlabeled data from the open domain can help adapt the federated model to new environments 78. Self-supervised learning techniques like contrastive learning can further facilitate this process 39.
  • Ensemble Learning: Combine multiple models trained on different clients via sophisticated ensemble methods can enhance the robustness of the final prediction, making it more generalizable to open domains 153.
  • Data Synthesis and Simulation: Generate synthetic data that mimics the characteristics of potential open domains can be used to pretrain or fine-tune the federated model, improving its generalization to unseen environments 33.
@@ 5.4.3 Burdens in Communication
Communication is a significant bottleneck in the implementation of FL in healthcare. In FL, each client needs to frequently communicate with the central server. This communication can be orders of magnitude slower than local computation due to constraints such as bandwidth, latency, and power 96.
The communication bottleneck in FL arises from several factors. First, the number of clients involved in an FL system can be very large (e.g., wearables and IoMT), leading to significant communication overhead. Each communication round requires sending model updates between the clients and the central server, which can be expensive in terms of time and resources 96, 4. Second, ensuring data privacy and security in FL is crucial, especially in sensitive domains like healthcare. The need for encryption and secure communication protocols adds to the computational and communication overhead. Encrypting model updates can significantly increase the size of the data being transmitted, further straining the communication channels and requiring more sophisticated algorithms to balance privacy and efficiency 4. Third, as FL tasks become more complex, the size of the models involved increases. Modern large-scale models, such as large language models (LLMs) and foundation models (FMs), can have billions of parameters, resulting in model sizes that require significant bandwidth to transmit. This issue is exacerbated when using standard communication protocols like gRPC, which have size limits on single messages (e.g., 2 GB). Typical LLMs and FMs can exceed these limits, necessitating the model to be split into smaller chunks for transmission, adding additional overhead and complexity to the system 166.
Included studies primarily concentrated on enhancing communication encryption techniques, with the aim of either reducing the volume of data exchanged 64, 105 or minimizing the number of communication rounds 56. Additionally, several studies explored methods for achieving fully decentralized communication without central server 57, 167. Improvements in encryption often involved methods for securely sharing secret keys among clients 84, 82, encryption mechanisms for safeguarding exchanged data 83, 65, and techniques for perturbing model outputs at each client using a secret key 168. To decrease the amount of data transmitted, some studies proposed transferring only a subset of model parameters 143, 169, 96 or employing strategies like compressing 98, masking 62, and quantizing gradients 96 or model outputs before exchange 27. Reducing the number of communication rounds was addressed through model design 92, 170, 99, aggregating updates based on elapsed time instead of epochs 98, 108, and evaluating the potential benefit of an update before communication 82. Other studies focused on detecting attacks during communication 171, developing authentication systems for clients 48, and improving client management systems 47, 65.
@@
  • Model Optimization: Utilize parameter-efficient, lightweight models such as MobileNet and SqueezeNet to reduce computation and memory overhead. Explore model pruning 172 and low-rank adaptation (LoRA) 173 techniques to decrease model size or regularize model weights without significantly impacting accuracy. Consider split learning 102, 25 to share intermediate activations or a subset of the model instead of the full model, thus reducing communication costs.
@@
  • Reduce Communication Rounds: Minimize the number of communication rounds by adopting methods such as One-Shot FL 174, which requires only a single round of communication to achieve effective model training.
@@
  • Knowledge Distillation: Use knowledge distillation techniques to transmit only essential distilled knowledge, such as logits from the final layer, rather than the entire model 154, 174, 52. This approach can reduce communication overhead while maintaining model performance.
  • Data Distillation: Apply data distillation to generate synthetic data summaries or distilled representations of the original datasets or updates. Clients can then share smaller, more concise updates with the central server, enhancing communication efficiency 175, 176.
@@ 5.4.4 Plain Convergence Analysis
Federated optimization in healthcare aims to adapt models to local data distributions while integrating global information. The inherent heterogeneity across hospitals and institutions often leads to instability and slow convergence in federated training 156. However, comprehensive convergence analysis is frequently lacking in current studies.
@@
Only a handful of studies 167, 62, 96, 105 have focused on the convergence of FL in healthcare settings. These studies typically relied on Stochastic Gradient Descent (SGD) as the foundational optimization method due to its effectiveness in smooth optimization problems, under assumptions such as the existence of lower bounds, Lipschitz smoothness, and bounded variance 167, 62. However, SGD-based FL algorithms often struggle with nonsmooth optimization problems, which are common in healthcare data due to irregularities in data distributions and the presence of outliers.
@@
  • Theoretical Convergence Guarantees: Establish rigorous theoretical methods that provide convergence guarantees for FL in healthcare. This includes deriving bounds on convergence rates and understanding the conditions under which proposed algorithms perform optimally, particularly in non-IID data settings. While complete boundary analyses in large-scale non-IID settings remain challenging due to data noise and complexity, partial modeling of noise, such as using bounded variance assumptions or noise-resilient gradient estimators, has shown promise in existing studies 156, 177. Future research could explore hybrid approaches that combine empirical noise estimation with theoretical bounds to enhance the practical relevance of convergence analysis.
@@
  • Advanced Optimization: Non-convexity, non-smoothness, and heterogeneity are not universal across all healthcare data. For instance, MRI images tend to be homogeneous, while blood test data exhibit more heterogeneity. A unified optimization method may not address all these challenges. Tailored optimization strategies, including smooth methods for homogeneous data and non-smooth, non-convex methods for heterogeneous data, should be explored. Hybrid approaches could also be considered to optimize convergence and stability in diverse healthcare applications.
@@ 5.4.5 Temporal Dynamics and Revoke Issues
Healthcare data’s inherent time dependence is critical, especially for diseases with distinct progression or treatment timelines, such as cancer and chronic conditions like diabetes. However, included studies often overlooked these temporal dynamics when partitioning data, potentially leading to models that inaccurately reflect disease progression. For example, COVID-19 characteristics, such as ground-glass opacities in lung CT scans, evolve with the disease’s progression 178. Ignoring such temporal dynamics can result in models that are overfitted to specific stages of a disease and unable to generalize across different phases 179, 180, 181. Furthermore, the dynamic nature of data at each participating hospital or institution complicates the situation. Hospitals continuously acquire new data and may also remove or modify existing data due to errors or other factors. This dynamic data environment requires FL models to adapt without frequent retraining, as new data might cause model drift, while data removal can leave critical gaps in the model’s understanding, particularly if the removed data represents rare or critical cases.
Another critical but overlooked issue in FL is data revocation, which becomes necessary when specific data must be withdrawn due to privacy concerns, regulatory requirements, patient or participant requests, misdiagnosis, invalidated prior diagnoses, or medical misconduct. Current FL setups, designed for iterative data aggregation, struggle with “unlearning” specific contributions without requiring complete model retraining. Emerging research highlights the need for mechanisms that allow for efficient data revocation without compromising the integrity of the model. For instance, methods have been proposed to facilitate client “unlearning” in FL, enabling the removal of data contributions from specific clients without significant degradation in model performance 182. This is critical in healthcare, where patient consent may be withdrawn, new privacy regulations may require data deletion, or misdiagnosed and fraudulent data could undermine model integrity.
@@
  • Data Revocation: Develop and integrate efficient data revocation techniques in FL frameworks. Approaches such as machine unlearning 182 can be refined to allow the removal of specific data contributions while minimizing the impact on overall model performance. This will be critical for maintaining compliance with privacy laws and upholding patient rights in healthcare applications.
@@ 5.4.6 Synchronization Issues
In Synchronous FL, all clients must complete their training and send their updates before the global model aggregation, which is straightforward but can be inefficient. This approach works well in environments where data is immediately available, such as in a centralized hospital picture archiving and communication system (PACS). However, in real-world scenarios, where data acquisition might be delayed due to network issues or the unavailability of input/output devices, synchronous updates can result in significant idle times and resource underutilization 61. Additionally, clients in an FL network, particularly smaller healthcare entities (e.g., wearables and IoMT devices), may not be active during every communication round, further delaying the global model update and potentially degrading overall system performance 4.
Asynchronous FL, on the other hand, allows clients to send updates independently, without waiting for other clients to finish their training. This approach is more flexible and can accommodate variations in client availability and computational power, thereby improving the overall efficiency of the FL process. One study 61 proposed an asynchronously updating FL architecture for cardiac activity monitoring and arrhythmia detection without the need for frequent synchronization. Another study 183 presented an adaptive asynchronous split FL scheme for medical image segmentation, which enhances training efficiency and model performance by allowing clients to operate at their own training speeds. However, asynchronous updates can introduce challenges related to the consistency of the global model, as updates from slower or less reliable clients may arrive out of sync with the rest of the system.
@@
Model Inversion 184 Reconstruct the actual samples (e.g., patient medical images or genetic data) from the model or updates Potential leakage of original patient data High Data privacy, patient confidentiality Anomaly detection in model outputs & updates DP
Membership Inference 185 Determining if a specific patient’s record was part of the model’s training set by analyzing its outputs Compromises patient confidentiality, leading to potential unauthorized access to sensitive health information Medium Data privacy, patient confidentiality Monitoring for unusual model behaviour, particularly overconfident predictions Regularization techniques, DP, and robust model validation
Data Attribute Inference 186 Infer individual sample’s attributes (like race, gender, age) or gain aggregate statistical insights about the entire training set from model parameters and updates Leakage of patient privacy Medium Data privacy, patient confidentiality, communication security Gradient analysis, privacy audits Secure aggregation, DP
Data Poisoning 187 Malicious participants alter local data or labels to degrade the global model’s performance Significant drop in model performance, potentially misleading patient diagnoses Medium Model accuracy, patient safety Monitoring for abnormal model behaviour, statistical checks Robust aggregation methods, anomaly detection in model updates
Model Poisoning 73 Malicious participants upload tampered model parameters to manipulate the global model Global model may be intentionally corrupted, affecting its performance on real data High Model integrity, patient safety Anomaly detection in model updates Robust aggregation, DP
Denial of Service 74 Disrupt the FL system by overwhelming it with requests or blocking normal data flow Training process delays or interruptions, affecting time-sensitive medical applications Medium System availability and security, training efficiency Monitoring network traffic for anomalies Rate limiting, robust network design, redundancy mechanisms
HE Attack 65 Decrypt encrypted model updates to access sensitive information Encryption key leakage may lead to a breakdown in data protection, compromising patient privacy High Data privacy, model integrity Encryption integrity checks, key management audits Secure key management, SMPC
User Withdrawal Attack 46 Participants withdraw from training, but their prior updates still affect the global model Updates from withdrawn users may contain errors or malicious data, degrading the global model Low Model integrity Tracking user participation and contribution consistency Periodic re-evaluation of model contributions, machine unlearning
@@ 5.5 Privacy and Security
Two critical privacy and security issues exist in current studies. First, it is concerning that 62% of the included studies did not encrypt model updates during communication. This lack of encryption leaves FL system vulnerable to interception, posing significant security risks. Second, statistical information such as sample sizes and distributions were often shared alongside model updates, particularly in FedAvg-based methods 107, 109. This practice can expose participants with large datasets to targeted attacks if an adversary intercepts the communication or compromises the aggregator 166.
Despite the advantage of FL in healthcare without directly exchanging or sharing local data, it is not immune to privacy and security risks. Adversaries can analyze changes in model updates over time to infer sensitive information or exploit system vulnerabilities to conduct targeted attacks using techniques like model inversion 188, membership inference 189, and poisoning 187. Additionally, clients may unintentionally reveal private data during the FL process. This can happen when the client memorizes previous model and gradient updates, leading to the leakage of sensitive information 66, 16. Furthermore, methods involving the sharing of a few data samples for augmentation or disclosing local data distributions during knowledge transfer can also result in privacy breaches 100, 52. These adversarial and unintentional exposures undermine the privacy and security of the FL process, necessitating robust countermeasures and continuous vigilance to safeguard the integrity and confidentiality of the FL process.
@@
Cryptography encompass a variety of methods, with homomorphic encryption (HE) 190 and secure multi-party computation (SMPC) 191 being among the most popular. HE enables computations directly on encrypted data without the need for decryption. This ensures that data remains encrypted throughout processing, storage, and transmission, thus providing robust data security and privacy. HE allows operations on ciphertexts, with the results, once decrypted, accurately reflecting the outcomes of operations performed on the original plaintext data. HE is particularly valuable in FL due to its ability to safeguard data privacy during computation. Recent research has highlighted HE’s effectiveness in various healthcare applications, including oncology and medical genetics. For instance, 70 demonstrated HE’s potential for truly private federated evaluations, and 78 successfully utilized HE for model aggregation in FL. While SMPC enables multiple parties to collaboratively compute a function over their combined data while keeping each party’s data private. Each participant holds a piece of encrypted or encoded data, and the computation is designed so that no party can access the others’ data or infer anything beyond the final result. SMPC ensures the confidentiality of both input data and computation results, making it robust against adversarial attacks and suitable for scenarios with multiple untrusted parties. It is increasingly used in healthcare and other sensitive applications to enhance privacy. For instance, research shows that SMPC can improve privacy in FL with medical datasets by addressing risks related to malicious models and enhancing the confidentiality of model aggregation 49, 73.
@@
Nevertheless, HE typically involves high storage and computational overheads. Encrypted data requires significant processing power, and the complexity of HE schemes can introduce a single point of failure, where a single server manages all encrypted data 70. Additionally, managing encryption keys securely is crucial, as any compromise in key management can jeopardize the entire system’s security 192. Also, SMPC often incurs high computational and communication overhead. The process of encrypting, encoding, and splitting data can be computationally intensive. Moreover, coordinating the computations across multiple parties requires substantial communication resources, which can become a bottleneck as the number of participants grows 193.
@@
Perturbation techniques, with differential privacy (DP) 194 being the most prominent, are crucial for protecting sensitive healthcare data. DP quantifies the risk of exposing individual data points and ensures that the inclusion or exclusion of any single data point has minimal impact on the model’s output by introducing randomness into the model’s results. In FL, DP is typically implemented by adding noise to the local updates or gradients before they are aggregated. This noise is designed to mask the contributions of individual data points, thus making it difficult to infer any single data point from the model’s outputs. DP provides a quantifiable measure of privacy through parameters such as $\epsilon$ and $\delta$, allowing for a clear understanding of the trade-off between privacy and model utility. Recent studies in healthcare have shown that FL with DP can achieve comparable accuracy to non-DP models in specific tasks, with a performance gap of less than 5%, proving DP’s efficacy with minimal accuracy compromise 107, 109, 166.
@@
However, the main challenge with DP is balancing the privacy budget with model performance. Adding noise typically reduces model accuracy by obscuring data patterns, which is a critical consideration in DP implementation 66. Managing the privacy budget $\epsilon$ involves a trade-off: a smaller $\epsilon$ enhances privacy but may reduce performance, and a larger $\epsilon$ offers less privacy protection. Effective management of this trade-off is essential, requiring careful attention to privacy needs and performance goals 130. Implementing DP also introduces additional computational overhead due to the noise addition and gradient clipping processes. This overhead can affect the scalability and efficiency of FL systems, particularly in scenarios with large numbers of clients or data 195.
@@
Beyond cryptography and perturbation methods, blockchain 23 and swarm learning (SL) 69, 24 have gained lots of attention for enhancing privacy and security in healthcare. Blockchain is increasingly being integrated with FL to address data security and trust issues by replacing the central server with a decentralized privacy protocol. The key advantage of blockchain lies in its distributed ledger system, which ensures that data is securely shared and maintained across all clients without relying on a central authority. This decentralization reduces the risk of single points of failure and enhances data integrity. Furthermore, blockchain’s verification schemes play a crucial role in the FL process, helping to detect and mitigate threats such as poisoning attacks. For instance, a blockchain-based FL framework developed by 76 combined DP and gradient-verification protocols to enhance security in IoMT devices, significantly reducing the success rate of poisoning attacks in tasks like diabetes prediction. Another approach by 75 utilized blockchain alongside an intrusion detection system to monitor and prevent malicious activities during model training, further securing federated healthcare networks. However, while blockchain offers robust security features, its integration with FL is often challenged by high communication and computational costs, which can limit its scalability and efficiency. Swarm Learning takes a different approach by decentralizing not just the privacy protocol but the entire model training process. In SL, there is no central aggregator, instead, decentralized hardware infrastructures work together to securely onboard clients and collaboratively generate a global model. This decentralized approach enhances the network’s resilience against attacks and is particularly effective in scenarios where data is non-IID, such as in the prediction of conditions like COVID-19 and leukemia 69, 24. However, SL’s reliance on peer-to-peer communication can introduce latency issues, which may slow down the training process, particularly in environments with varying network conditions.